OpenStudy (anonymous):

PLEASE HELP! I WILL MEDAL! Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). The possible answers are: a. -1/3x-6 b. -1/3x-14 c. 3x-6 d. 3x-14

OpenStudy (anonymous):

Firstly, if it's a parallel line what should be its slope ?

OpenStudy (anonymous):

Do you how to find?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okey we can write lines either two forms: ax^2+bx+c or y = mx+n In first form the slope will be -a/b and in second form m is the slope. If we organize the equation we get 3x-y=-3 3x=y+3 y=3x-3 So slope should be 3.

OpenStudy (anonymous):

Sorry it shoul be 3x=y-3 in second step Then y=3x+3

OpenStudy (anonymous):

so how do you find the answer to the problem?

OpenStudy (anonymous):

Okey moving on we found that the slope should be 3. So as I said before you can write a line like y=3x+n because we know the slope but we don't know what exact line is. It says line passes through the point (4,-2) It should be true for our line 4 is x and -2 is y (x,y) -2=3*4+n -2=12+n n=-14 The line is y=3x-14